# A nonperturbative test of nucleation calculations for strong phase transitions

### v2

@@ I High-temperature effective field theory
At high temperatures, the equilibrium thermodynamics of a weakly coupled 3+1 dimensional quantum field theory is captured by a 3 dimensional effective field theory containing only the light bosonic fields [71, 72]. The parameters of this effective theory depend on those of the full theory and on the temperature. The precise matching relations can be computed order-by-order in powers of couplings.
In the vicinity of a phase transition, the effective mass of the field undergoing the transition becomes small, and hence the longest wavelength modes may often be described by an effective theory for this field alone, with all other fields integrated out. We will assume that this is the case for the real scalar field, in which case the effective 3d Lagrangian is
@@
At the critical temperature (in this case $T_{\text{c}}=98.513$ GeV), this goes through the second benchmark point studied in Ref. [61]. The critical temperature is exact within the 3d EFT, as it is protected by a $Z_{2}$ symmetry. The equilibrium thermodynamics at the critical temperature was found to be under good perturbative control at one-loop order, at which point the dimensionless loop-expansion parameter within the EFT is $\alpha_{3}\approx 0.13$ [61]. The temperature of spinodal decomposition, where the mass in the metastable phase goes through zero, is $T_{\text{s}}=89.920$ GeV at tree-level, though this is corrected by loops within the EFT.
Our lattice simulations are carried out within the 3d effective theory. The majority of our simulations were carried out at the parameter point corresponding to $T=93.121$ GeV, where we have used Eqs. (18) to (21) to fix the corresponding 3d effective parameters. The dimensionless loop-expansion parameter at this temperature is $\alpha_{3}\approx 0.23$ [61]. Note that, while we have used only leading order dimensional reduction matching relations, our calculations within the 3d effective theory are valid nonperturbatively at the values of the effective parameters studied.
@@ III Lattice details
Here we give further details of our lattice discretisation. We outline the purely spatial action used in the Monte-Carlo simulations, as well as details of our real-time update algorithm. Our simulation code is available at https://bitbucket.org/og113/scalnuc/.
@@
$\pi_{t+\frac{1}{2},x}$ $=\pi_{t,x}-\frac{\partial H_{\text{eff}}}{\partial\phi_{t,x}}\frac{\Delta t}{2},$ (37)
$\phi_{t+1,x}$ $=\phi_{t,x}+a^{3}\pi_{t+\frac{1}{2},x}\Delta t,$ (38)
$\pi_{t+1,x}$ $=\pi_{t+\frac{1}{2},x}-\frac{\partial H_{\text{eff}}}{\partial\phi_{t+1,x}}\frac{\Delta t}{2}.$ (39)
@@
$\vartheta$ $=1-\exp(-2a^{3}\gamma\Delta t).$ (44)

### v3

@@ I High-temperature effective field theory
At high temperatures, the equilibrium thermodynamics of a weakly coupled 3+1 dimensional quantum field theory is captured by a 3 dimensional EFT containing only the light bosonic fields [59, 60]. The parameters of this EFT depend on those of the full theory and on the temperature. The precise matching relations can be computed order-by-order in powers of couplings.
In the vicinity of a phase transition, the effective mass of the field undergoing the transition becomes small, and hence the longest wavelength modes may often be described by an EFT for this field alone, with all other fields integrated out. We will assume that this is the case for the real scalar field, in which case the effective 3d Lagrangian is
@@
At the critical temperature (in this case $T_{\text{c}}=98.513~\text{GeV}$), this goes through the second benchmark point studied in Ref. [51]. The critical temperature is exact within the 3d EFT, as it is protected by a $Z_{2}$ symmetry. The equilibrium thermodynamics at the critical temperature was found to be under good perturbative control at one-loop order, at which point the dimensionless loop-expansion parameter within the EFT is $\alpha_{3}\approx 0.13$ [51]. The temperature of spinodal decomposition, where the mass in the metastable phase goes through zero, is $T_{\text{s}}=89.920~\text{GeV}$ at tree-level, though this is corrected by loops within the EFT.
Our lattice simulations are carried out within the 3d EFT. The majority of our simulations were carried out at the parameter point corresponding to $T=93.121~\text{GeV}$, where we have used Eqs. (20) to (23) to fix the corresponding 3d effective parameters. The dimensionless loop-expansion parameter at this temperature is $\alpha_{3}\approx 0.23$ [51]. Note that, while we have used only leading order dimensional reduction matching relations, our calculations within the 3d EFT are valid nonperturbatively at the values of the effective parameters studied.
@@ III Lattice details
Here we give further details of our lattice discretisation. We outline the purely spatial action used in the Monte-Carlo simulations, as well as details of our real-time update algorithm. Our simulation code is available at [74].
@@
$\pi_{t+\frac{1}{2},x}$ $=\pi_{t,x}-\frac{1}{a^{3}}\frac{\partial H_{\text{eff}}}{\partial\phi_{t,x}}\frac{\Delta t}{2},$ (39)
$\phi_{t+1,x}$ $=\phi_{t,x}+\pi_{t+\frac{1}{2},x}\Delta t,$ (40)
$\pi_{t+1,x}$ $=\pi_{t+\frac{1}{2},x}-\frac{1}{a^{3}}\frac{\partial H_{\text{eff}}}{\partial\phi_{t+1,x}}\frac{\Delta t}{2}.$ (41)
@@
$\vartheta^{2}$ $=1-\exp(-2\gamma\Delta t).$ (46)